# NL-Debugging: Exploiting Natural Language as an Intermediate Representation for Code Debugging

### v1

@@ Abstract
Weiming Zhang1thanks: Equal Contribution., Qingyao Li [^1] footnotemark: 1, Xinyi Dai2, Jizheng Chen1, Kounianhua Du1, Weinan Zhang1thanks: Corresponding author., Weiwen Liu1, Yasheng Wang2, Ruiming Tang2, Yong Yu1 1Shanghai Jiao Tong University, 2Huawei Noah’s Ark Lab Shanghai, China {WeimingZhang_2020, wnzhang}@sjtu.edu.cn
@@ 1 Introduction
Code generation is a challenging task that requires strong reasoning abilities and a deep understanding of programming languages. While large language models (LLMs) demonstrate potential in this domain Achiam et al. (2023); Zhao et al. (2023); Touvron et al. (2023); Zhu et al. (2024); Li et al. (2023), they often struggle to produce fully correct code implementations in one attempt Liu et al. (2024); Dou et al. (2024). This limitation has driven recent research toward iterative refinement approaches for code generation Chen et al. (2023); Zhong et al. (2024), where the debugging process—transforming flawed code into correct implementations—has emerged as a critical research focus Jin et al. (2023).
Early debugging methodologies primarily relied on execution feedback for code-level analysis Zhang et al. (2023a); Chen et al. (2023); Zhong et al. (2024). Although effective for identifying syntax errors or fundamental logical flaws, these approaches show limited capability in detecting deeper algorithmic design issues Finder and Fey (2010); Tian et al. (2024). Inspired by the success of natural language reasoning in LLMs Jaech et al. (2024); Guo et al. (2025); Qin et al. (2024), recent advancements have shifted focus to natural language-based code optimization Wang et al. (2024); Li et al. (2024b, a). This trend reflects a growing recognition of the potential for natural language to serve as a powerful medium for guiding and improving code-related tasks.
@@ 2.2 Code Generation with Natural Language Reasoning
Recent advances in LLM reasoning have demonstrated strong potential in handling complex tasks, with natural language playing a central role Huang and Chang (2022). Initially, Chain-of-Thought (CoT) methods Wei et al. (2022); Wang et al. (2022) organized reasoning into step-by-step natural language explanations. Building on this, plan-based approaches like Tree-of-Thought (ToT) Yao et al. (2024) and Graph-of-Thought Besta et al. (2024) introduced structured search spaces to explore multiple reasoning paths, expanding the solution space. More recent methods such as o1 Jaech et al. (2024), STaR Zelikman et al. (2022), and Iter-MCTS Xie et al. (2024) focus on iterative refinement, enabling models to progressively improve their reasoning.
Recent research has started incorporating natural language reasoning explicitly into code generation. For instance, explorations like RethinkMCTS Li et al. (2024b) and CodeTree Li et al. (2024a) employ tree search techniques within the natural language space, iteratively refining reasoning steps to improve code quality. However, many existing methods assume a fixed natural language format and overlook how different structures can impact reasoning effectiveness. This motivates our investigation into how varying natural language representations can better support LLM-driven code generation and debugging.
@@ 4.1.1 Datasets
We evaluate our NL-Debugging framework on two widely used code datasets: APPS Hendrycks et al. (2021) and Codeforces (MatrixStudio, 2024). The APPS dataset consists of three difficulty levels—introductory, interview, and competition. From each difficulty level, we selected 100 problems to ensure a balanced assessment. The Codeforces dataset includes problems from online programming contests categorized by "ratings". We chose problems with ratings of 1200, 1500, and 1800 and selected 100 problems per rating level. Within each difficulty and rating category, problems were randomly selected to ensure balanced coverage, following Zhang et al. (2023b); Islam et al. (2024). We choose the first 100 problems per difficulty or rating level to maintain randomness.
@@ 4.1.2 Metrics
GPT4o-mini Base(w/o debugging) 56.45 54.57 34.67 35 28 16 64.06 47.60 36.60 42 23 15
@@ 4.1.4 Implementation
We select GPT-4o-mini OpenAI (2024), Claude-3.5-sonnet Anthoropic (2024), and DeepSeek-Coder-V2-Lite Zhu et al. (2024) as the LLM backbones for our experiments. GPT-4o-mini and Claude-3.5-sonnet are accessed via their respective API forms, while DeepSeek-Coder-V2-Lite (16B) is locally deployed. Since all the methods utilize an iterative debugging process, we set the maximum debugging time to 5 iterations. All models are run with a temperature of 0. The source code is available for reviewers  [^1] Source code: https://anonymous.4open.science/r/NL-Debugging-4280/ and will be publicly released upon acceptance.
@@ 4.2 Code Debugging Performance (RQ1)
  • At the APPS Introductory level, DeepSeek-Coder-V2-Lite shows minimal improvement in pass@1 compared to baseline methods. While its pass rate surpasses some baselines, its debugging gains are less significant than those of GPT4o-mini and Claude3.5-Sonnet. This likely stems from DeepSeek-Coder-V2-Lite’s smaller parameter size and coder-focused design, which may limit its capacity for complex reasoning.
@@ 4.3 What Format of Natural Language Representation Works the Best? (RQ2)
  • Key points: This format abstracts the code’s main logic into concise key points, summarizing the key thought steps and operations Li et al. (2024b). Key points are iteratively added and refined during debugging to provide a simplified, high-level view of the algorithm.
@@ 4.5 How does Natural Language Work Better with Code Execution Feedback? (RQ4)
NL-Debugging 71.36 69.96 44.17 51 48 23 71.36 69.96 44.17 51 48 23
Long CoT(Direct) 67.05 63.73 42.50 48 40 23 67.05 63.73 42.50 48 40 23
Long CoT(Sketch) 71.08 68.16 45.17 51 47 26 71.08 68.16 45.17 51 47 26
@@ Appendix A Dataset Details
We evaluate our NL-Debugging framework on two widely used code datasets: APPS Hendrycks et al. (2021) and Codeforces (MatrixStudio, 2024). The APPS dataset includes three difficulty levels—introductory, interview, and competition—with a total of 5000 programming problems for training and 5000 for testing. The Codeforces dataset contains problems from the Codeforces online programming contest, with varying difficulty levels categorized by "ratings", from which we choose 1200, 1500, and 1800 for evaluation.
@@ Appendix C Metric Details
Appendix D More Details on Experimental Settings
D.1 Code Debugging vs. Code Generation: Their Connection
@@ Appendix D More Details on Experimental Settings
Figure 6: A case study for NL-Debugging.
D.2 Code Debugging vs. Automatic Program Repair: Their Difference
@@
Appendix E More Experimental Results
The experimental findings yield the following conclusions:
Appendix F Case Study
@@ G.1 Prompts for NL-Debugging
Prompt for Backtranslation You are an expert Python programmer. Below is an algorithmic question (problem specification) along with the current implementation for solving the problem. $\{\text{Problem Description}\}{ Problem Description }$ $\{\text{Current Code Implementation}\}{ Current Code Implementation }$ Your task is to generate a **Natural Language Sketch** for this code. This sketch should describe the logical reasoning or steps that the code is trying to follow in order to solve the problem. Do not focus on syntax or specific code lines, but explain the thought process or approach the code takes to solve the problem at a high level.
@@
[…]
@@
[…]
@@
Prompt for Regeneration You are an expert Python programmer. Below is a high-level natural language sketch for a correct solution to an algorithmic problem, along with the problem description. $\{\text{Problem Description}\}{ Problem Description }$ $\{\text{Refined Natural Language Sketch}\}{ Refined Natural Language Sketch }$ Your task is to write Python code that implements this sketch and solves the problem. Do not include unnecessary comments or explanations, only the code itself.
@@ G.2 Prompts for Different Types of Natural Language
Prompt for Key Points Backtranslation You are an expert Python programmer. Below is an algorithmic question (problem specification) along with the current implementation for solving the problem. $\{\text{Problem Description}\}{ Problem Description }$ $\{\text{Current Code Implementation}\}{ Current Code Implementation }$ I need you to extract 3 key points (thoughts) that summarize the core algorithm or logic used in this code. Please list each thought in a separate entry in the format: The output should be a **list of dicts** with each key as ‘Thought-i‘. Do not include explanations or justifications, just focus on capturing the key algorithmic points. [ {"Thought-1": "We could use the print function to finish the task in one line: print(2 + 3)."}, {"Thought-2": "We should calculate the problem by setting a=2+3, and then print(a)."}, {"Thought-3": "The problem can’t be solved by Python."} ]
@@
Prompt for Key Points Refinement You are an expert Python programmer. Below is an algorithmic problem description, the current list of thoughts (which contains bugs), the current code implementation, and the feedback from running the code. $\{\text{Problem Description}\}{ Problem Description }$ $\{\text{Key Points with Bugs}\}{ Key Points with Bugs }$ $\{\text{Current Code Implementation}\}{ Current Code Implementation }$ $\{\text{Bug Analysis}\}{ Bug Analysis }$ Please analyze and generate 1-2 new thoughts to correct the approach. **Format Requirements:** 1. Output ONLY a JSON-parsable list of dictionaries 2. Each dictionary must use ’Thought-i’ as key (i continues from previous sequence) 3. Each value should concisely state one algorithmic insight. 3. New thoughts can correct the previous buggy thoughts. [ {"Thought-4": "Using depth-first search for tree traversal"}, {"Thought-5": "Handling leaf nodes with null checks"}, ]
@@
Prompt for Pseudocode Backtranslation You are an expert Python programmer. Below is an algorithmic question (problem specification) along with the current implementation for solving the problem. $\{\text{Problem Description}\}{ Problem Description }$ $\{\text{Current Code Implementation}\}{ Current Code Implementation }$ Your task is to generate **pseudocode** for this code in LaTeX format using the algorithm2e package. This pseudocode should provide the high-level logic of the code. Format it as LaTeX pseudocode that captures the main steps and logic of the code.
@@
Prompt for Pseudocode Refinement You are an expert Python programmer. Below is an algorithmic problem description, the current pseudocode of the solution (which contains bugs), the current code implementation, and the feedback from running the code. $\{\text{Problem Description}\}{ Problem Description }$ $\{\text{Pseudocode with Bugs}\}{ Pseudocode with Bugs }$ $\{\text{Current Code Implementation}\}{ Current Code Implementation }$ $\{\text{Bug Analysis}\}{ Bug Analysis }$ Please rethink the approach and generate a new, corrected pseudocode for solving the problem referring to the bug explanation for this buggy pseudocode. This new pseudocode should address the issues in the previous pseudocode, focus on outlining the corrected steps needed to solve the problem, and avoid the errors found in the previous pseudocode."

### v2

@@ Abstract
Weiming Zhang1thanks: Equal Contribution., Qingyao Li [^1] footnotemark: 1, Xinyi Dai1, Jizheng Chen1, Kounianhua Du1, Weiwen Liu1thanks: Corresponding author., Yasheng Wang2, Ruiming Tang2, Yong Yu1, Weinan Zhang1 1Shanghai Jiao Tong University, 2Huawei Noah’s Ark Lab Shanghai, China {WeimingZhang_2020, ly890306, wwliu}@sjtu.edu.cn
@@ 1 Introduction
Code generation is a challenging task that requires strong reasoning abilities and a deep understanding of programming languages. While large language models (LLMs) demonstrate potential in this domain Achiam et al. (2023); Zhu et al. (2024); Li et al. (2023), they often struggle to produce fully correct code implementations in one attempt Liu et al. (2024); Dou et al. (2024). This limitation has driven recent research toward iterative refinement approaches for code generation Chen et al. (2023); Zhong et al. (2024); Chen et al. (2025), where the debugging process—transforming flawed code into correct implementations—has emerged as a critical research focus.
Early debugging methodologies primarily relied on execution feedback for code-level analysis Zhang et al. (2023a); Chen et al. (2023); Zhong et al. (2024). Although effective for identifying syntax errors or fundamental logical flaws, these approaches show limited capability in detecting deeper algorithmic design issues Finder and Fey (2010); Tian et al. (2024). Inspired by the success of natural language reasoning in LLMs Jaech et al. (2024); Guo et al. (2025); Qin et al. (2024), recent advancements have shifted focus to natural language-based code optimization Wang et al. (2024); Li et al. (2024, 2025a). This trend reflects a growing recognition of the potential for natural language to serve as a powerful medium for guiding and improving code-related tasks.
@@ 2.2 Code Generation with Natural Language Reasoning
Recent advances in LLM reasoning have demonstrated strong potential in handling complex tasks, with natural language playing a central role Huang and Chang (2022); Yang et al. (2025). Initially, Chain-of-Thought (CoT) methods Wei et al. (2022); Wang et al. (2022) organized reasoning into step-by-step natural language explanations. Building on this, plan-based approaches like Tree-of-Thought (ToT) Yao et al. (2024) and Graph-of-Thought Besta et al. (2024) introduced structured search spaces to explore multiple reasoning paths, expanding the solution space. More recent methods such as o1 Jaech et al. (2024), STaR Zelikman et al. (2022), and Iter-MCTS Xie et al. (2024) focus on iterative refinement, enabling models to progressively improve reasoning Li et al. (2025b); Zhu et al. (2025).
Recent research has started incorporating natural language reasoning explicitly into code generation. For instance, explorations like RethinkMCTS Li et al. (2024) and CodeTree Li et al. (2025a) employ tree search techniques within the natural language space, iteratively refining reasoning steps to improve code quality. However, many existing methods assume a fixed natural language format and overlook how different structures can impact reasoning effectiveness. This motivates our investigation into how varying natural language representations can better support LLM-driven code generation and debugging.
@@ 4.1.1 Datasets
We evaluate our NL-Debugging framework on two widely used code datasets: APPS Hendrycks et al. (2021) and Codeforces (Team, 2024). The APPS dataset consists of three difficulty levels—introductory, interview, and competition. From each difficulty level, we selected 100 problems to ensure a balanced assessment. The Codeforces dataset includes problems from online programming contests categorized by "ratings". We chose problems with ratings of 1200, 1500, and 1800 and selected 100 problems per rating level. Within each difficulty and rating category, problems were randomly selected to ensure balanced coverage, following Zhang et al. (2023b); Islam et al. (2024). We choose the first 100 problems per difficulty or rating level to maintain randomness.
@@ 4.1.2 Metrics
GPT-4o-mini Base(w/o debugging) 56.45 54.57 34.67 35 28 16 64.06 47.60 36.60 42 23 15
@@ 4.1.4 Implementation
We select GPT-4o-mini OpenAI (2024), Claude-3.5-sonnet Anthoropic (2024), and DeepSeek-Coder-V2-Lite Zhu et al. (2024) as the LLM backbones for our experiments. GPT-4o-mini and Claude-3.5-sonnet are accessed via their respective API forms, while DeepSeek-Coder-V2-Lite (16B) is locally deployed. Since all the methods utilize an iterative debugging process, we set the maximum debugging time to 5 iterations. All models are run with a temperature of 0.  [^1] The source code of this work is made available at https://github.com/yevzh/NL-Debugging.
@@ 4.2 Code Debugging Performance (RQ1)
  • At the APPS Introductory level, DeepSeek-Coder-V2-Lite shows minimal improvement in pass@1 compared to baseline methods. While its pass rate surpasses some baselines, its debugging gains are less significant than those of GPT-4o-mini and Claude3.5-Sonnet. This likely stems from DeepSeek-Coder-V2-Lite’s smaller parameter size and coder-focused design, which may limit its capacity for complex reasoning.
@@ 4.3 What Format of Natural Language Representation Works the Best? (RQ2)
  • Key points: This format abstracts the code’s main logic into concise key points, summarizing the key thought steps and operations Li et al. (2024). Key points are iteratively added and refined during debugging to provide a simplified, high-level view of the algorithm.
@@ 4.5 How does Natural Language Work Better with Code Execution Feedback? (RQ4)
NL-Debugging 71.36 69.96 44.17 51 48 23 79.57 64.00 48.69 65 41 25
Long CoT(Direct) 67.05 63.73 42.50 48 40 23 74.50 62.14 43.79 55 37 21
Long CoT(Sketch) 71.08 68.16 45.17 51 47 26 77.36 63.96 45.12 60 38 22
@@ 5 Conclusion
Acknowledgements
The Shanghai Jiao Tong University team is partially supported by National Key R&D Program of China (2022ZD0114804), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and National Natural Science Foundation of China (62322603, 62177033, 62502310).
@@ Appendix A Dataset Details
We evaluate our NL-Debugging framework on two widely used code datasets: APPS Hendrycks et al. (2021) and Codeforces (Team, 2024). The APPS dataset includes three difficulty levels—introductory, interview, and competition—with a total of 5000 programming problems for training and 5000 for testing. The Codeforces dataset contains problems from the Codeforces online programming contest, with varying difficulty levels categorized by "ratings", from which we choose 1200, 1500, and 1800 for evaluation.
@@ Appendix C Metric Details
Figure 6: A case study for NL-Debugging.
Appendix D More Experimental Results
The experimental findings yield the following conclusions:
Appendix E Case Study
@@
Appendix F More Details on Experimental Settings
F.1 Code Debugging vs. Code Generation: Their Connection
F.2 Code Debugging vs. Automatic Program Repair: Their Difference
@@ G.1 Prompts for NL-Debugging
Prompt for Backtranslation You are an expert Python programmer. Below is an algorithmic question (problem specification) along with the current implementation for solving the problem. $\{\text{Problem Description}\}$ $\{\text{Current Code Implementation}\}$ Your task is to generate a **Natural Language Sketch** for this code. This sketch should describe the logical reasoning or steps that the code is trying to follow in order to solve the problem. Do not focus on syntax or specific code lines, but explain the thought process or approach the code takes to solve the problem at a high level.
@@
Prompt for Execution Analysis You are an expert Python programmer. You will be provided with an algorithmic problem description, the current Python code implementation, and the execution feedback that indicates where the code went wrong. $\{\text{Problem Description}\}$ $\{\text{Current Code Implementation}\}$ $\{\text{Execution Feedback}\}$ $\{\text{Natural Language Sketch with Bugs}\}$ Please analyze the feedback and provide an explanation of what went wrong in the code and why it failed in this sketch. Do not provide specific steps to fix the sketch. Focus solely on explaining the root cause of the issue in two or three sentences.
@@
Prompt for Natural Language Refinement You are an expert Python programmer. Below is an algorithmic problem description, the current natural language sketch of the solution (which contains bugs), the current code implementation, and the feedback from running the code, as well as a detailed expert analysis of the bug. $\{\text{Problem Description}\}$ $\{\text{Natural Language Sketch with Bugs}\}$ $\{\text{Current Code Implementation}\}$ $\{\text{Bug Analysis}\}$ Based on the feedback and expert analysis of the current sketch, please provide a refined and corrected version of the sketch. The corrected sketch should: 1. Identify and correct specific points in the current sketch where errors or incorrect assumptions have been made. 2. Expand and elaborate on problematic steps in greater detail to explain the correct reasoning, ensuring that you address and fix the issues identified in the previous sketch. 3. Ensure that each step is logically connected and that any potential issues in the original approach are explicitly avoided in the new version. Do not include any specific code. The goal is to refine and improve the high-level natural language explanation of the problem-solving approach.
@@
Prompt for Regeneration You are an expert Python programmer. Below is a high-level natural language sketch for a correct solution to an algorithmic problem, along with the problem description. $\{\text{Problem Description}\}$ $\{\text{Refined Natural Language Sketch}\}$ Your task is to write Python code that implements this sketch and solves the problem. Do not include unnecessary comments or explanations, only the code itself.
@@ G.2 Prompts for Different Types of Natural Language
Prompt for Key Points Backtranslation You are an expert Python programmer. Below is an algorithmic question (problem specification) along with the current implementation for solving the problem. $\{\text{Problem Description}\}$ $\{\text{Current Code Implementation}\}$ I need you to extract 3 key points (thoughts) that summarize the core algorithm or logic used in this code. Please list each thought in a separate entry in the format: The output should be a **list of dicts** with each key as ‘Thought-i‘. Do not include explanations or justifications, just focus on capturing the key algorithmic points. [ {"Thought-1": "We could use the print function to finish the task in one line: print(2 + 3)."}, {"Thought-2": "We should calculate the problem by setting a=2+3, and then print(a)."}, {"Thought-3": "The problem can’t be solved by Python."} ]
@@
Prompt for Key Points Refinement You are an expert Python programmer. Below is an algorithmic problem description, the current list of thoughts (which contains bugs), the current code implementation, and the feedback from running the code. $\{\text{Problem Description}\}$ $\{\text{Key Points with Bugs}\}$ $\{\text{Current Code Implementation}\}$ $\{\text{Bug Analysis}\}$ Please analyze and generate 1-2 new thoughts to correct the approach. **Format Requirements:** 1. Output ONLY a JSON-parsable list of dictionaries 2. Each dictionary must use ’Thought-i’ as key (i continues from previous sequence) 3. Each value should concisely state one algorithmic insight. 3. New thoughts can correct the previous buggy thoughts. [ {"Thought-4": "Using depth-first search for tree traversal"}, {"Thought-5": "Handling leaf nodes with null checks"}, ]
@@
Prompt for Pseudocode Backtranslation You are an expert Python programmer. Below is an algorithmic question (problem specification) along with the current implementation for solving the problem. $\{\text{Problem Description}\}$ $\{\text{Current Code Implementation}\}$ Your task is to generate **pseudocode** for this code in LaTeX format using the algorithm2e package. This pseudocode should provide the high-level logic of the code. Format it as LaTeX pseudocode that captures the main steps and logic of the code.
@@
Prompt for Pseudocode Refinement You are an expert Python programmer. Below is an algorithmic problem description, the current pseudocode of the solution (which contains bugs), the current code implementation, and the feedback from running the code. $\{\text{Problem Description}\}$ $\{\text{Pseudocode with Bugs}\}$ $\{\text{Current Code Implementation}\}$ $\{\text{Bug Analysis}\}$ Please rethink the approach and generate a new, corrected pseudocode for solving the problem referring to the bug explanation for this buggy pseudocode. This new pseudocode should address the issues in the previous pseudocode, focus on outlining the corrected steps needed to solve the problem, and avoid the errors found in the previous pseudocode."